OpenStudy (princessashley):

What are the roots of 2x+6= -5/x?

OpenStudy (mww):

As the root is not x = 0, multiply both sides by x and solve the quadratic you result in.

OpenStudy (pythagoras123):

Even if the root was x=0, you are still allowed to multiply on both sides by x. It's dividing on both sides by x that you should avoid. \[2x ^{2}+6x=-5\] \[2x ^{2}+6x+5= 0\] \[x=?\]

OpenStudy (luhvon):

try one of these and see if it works x =(2-√76)/6=(1-√ 19 )/3= -1.120 2. x =(2+√76)/6=(1+√ 19 )/3= 1.786

jhonyy9 (jhonyy9):

2x+6 = -5/x interdiction of x that not can being equal by zero bc. than this fraction of -5/x will be undefined 2x+6=-5/x multiple both sides by x 2x^2 +6x = -5 2x^2 +6x +5 = 0 D = 36 -40 = - 4 x_1,2 = (-6 +/- sqrt(-4)/4 = (-6 +/- 2i)/4 = (-3+/- i)/2 hope this will help you